Circular No. 8070

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ao IN NGC 2993 Further to IAUC 8069, J. Graham and W. Li report the LOTOSS

discovery of apparent supernova (mag ∼ 17.3) on an unfiltered KAIT image taken on Feb. 10.4 UT. The new object is confirmed at about the same brightness in an earlier image taken on Jan. 30.4. SN 2003ao is located at α = 9^{h}45^{m}48^{s}.46, δ = −14^{o}22′14′′.0 (equinox 2000.0), which is 2′′.3 east and 8′′.4 south of the nucleus of NGC 2993. A KAIT image taken on 2002 Dec. 2.4 showed nothing at this position (limiting mag ∼ 19.0).

X1901+031 D. Galloway, R. Remillard, and E. Morgan, Massachusetts Institute

of Technology; and J. Swank, Goddard Space Flight Center, on behalf of the RXTE Instrument Teams, report a new x-ray outburst in a transient last seen in 1970–1971 (Forman et al. 1976, Ap.J. 206, L29): “The ASM position is α = 19^{h}03^{m}33^{s}, δ = +3^{o}12′.4 (equinox J2000.0; estimated 3′ uncertainty at 90-percent confidence). This position is within the 10′ error box originally estimated for X1901+031. The current outburst appears to have begun on 2003 Jan. 31 with a flux of 40(8) mCrab at 2–10 keV, rising to a steady flux of 130 mCrab on Feb. 9 and 10. Brief follow-up observations with the RXTE PCA instrument on Feb. 10 show pulsations with a period of 2.763 s. We encourage efforts to locate the optical counterpart, noting a likely optical companion with an early-type spectrum or Be classification.”

IGR J16318–4848 H. Murakami and T. Dotani, Institute of Space and Astronautical Sci-

ence; and R. Wijnands, University of St Andrews, report on the ASCA archival data analysis of the field of IGR J16318–4848 (IAUC 8063). A clear x-ray source was detected with both GIS and SIS at a position consistent with that of IGR J16318–4848, from a net exposure of 4000 s on 1994 Sept. 3.2 UT. The SIS position was determined as α = 16^{h}31^{m}49^{s}, δ = −48^{o}49′.2 (equinox 2000.0; uncertainty ∼ 0′.8). Murakami et al. add: “The source had a highly absorbed spectrum. If we assume a power-law continuum with a photon index fixed to 1.0, the neutral-hydrogen-absorption column density is estimated to be 4 ± 2 × 10^{23} cm^{−2}. The source flux was ∼ 4 × 10^{−11} erg s^{−1} cm^{−2} in 2–10 keV when corrected for the absorption. Although the data statistics were not very high, we noticed a hint of an iron emission line at 6.4 keV.”

2003 February 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green